# SpecFuse: Ensembling Large Language Models via Next-Segment Prediction

### v1

@@ 6 Conclusion
In this paper, we introduce SpecFuse, a novel ensemble framework that generates fused outputs by iteratively producing the next segment through collaboration among LLMs, allowing base LLMs to be seamlessly integrated without any training or adaptation. Additionally, SpecFuse employs a model exit mechanism that dynamically excludes underperforming models in previous rounds during query responses, reducing computational costs. Experimental results across six benchmarks demonstrate that SpecFuse consistently delivers more stable performance compared to single LLMs and previous ensemble methods. We hope our work inspires further research on online model ensemble, improving the quality of responses delivered to users based on existing LLMs.
Appendix A Dataset Details
The following provides a detailed description of the evaluation of instruction-response capability using the datasets.
@@ Appendix A Dataset Details
For the English dataset, we choose the Alpaca-gpt4 (Peng et al., 2023) and Dolly-15k (Conover et al., 2023) datasets for evaluation, both of which have inputs that consist of human instructions. We select these two datasets because their response sources differ: the Dolly-15k dataset features human-provided responses, while the Alpaca-GPT-4 dataset contains responses generated by the state-of-the-art GPT-4 (OpenAI et al., 2024) model, which provides neutral answers to each question and can refuse to answer inappropriate or harmful ones. Using both types of responses for scoring allows us to more thoroughly compare the advantages of our ensemble system. Additionally, due to the large size of these datasets, we randomly sample portions from each to create a new test set. From the Dolly-15k dataset, we randomly select 1,500 open-QA samples for testing, with 500 reserved for the development set. In the Alpaca-GPT-4 dataset, after shuffling the data, we manually verify the correctness of GPT-4’s responses and select 2,000 validated samples, with 1,500 used for testing and 500 for validation.
For the Chinese dataset, we utilize the Human-Value and Ruozb datasets from the COIG-CQIA (Bai et al., 2024) benchmark for testing. The instructions in these two datasets consist of human-posed questions, with answers provided either by humans or generated by GPT-4. The COIA authors manually review and filter the responses, retaining only the correct answers generated by GPT-4.
Appendix B Evaluation Methods
To evaluate the quality of our framework’s responses to human questions in the dataset, a range of metrics assessing model generation capabilities are selected for the following experiments.
  • $\bullet$
@@ Appendix B Evaluation Methods
  • $\bullet$
    The GPT4-Rank evaluation utilizes the GPT-$4$The version we use is GPT-4-turbo, and the link is https://openai.com/index/gpt-4/ model to compare two different responses against a ground-truth response. The model will select the better of the two responses. For each test sample, we pair the responses generated by different models and have GPT-4 determine which one is superior. Since the MBR and PairRank methods do not generate new responses, we do not re-rank the responses they select from the base LLMs. Instead, we use the average rankings of the responses they select from the base LLMs to represent their GPT4-Rank. Once all comparisons are complete, we count the number of wins for each model. Based on these win counts, we rank the responses from the different models. The average ranking of each model across all data in the dataset is the value reported in our table. The evaluation instructions for GPT-4 are shown in Table 5.
@@ 2. B is better
Appendix C Case Study

### v2

@@ 5 Conclusion
In this paper, we introduce SpecFuse, a novel ensemble framework that generates fused outputs by iteratively producing the next segment through collaboration among LLMs, allowing base LLMs to be seamlessly integrated without any training or adaptation. Additionally, SpecFuse employs a model exit mechanism that dynamically excludes underperforming models in previous rounds during query responses, reducing computational costs. We hope our work inspires further research on model ensemble, improving the quality of responses delivered to users based on existing LLMs.
Limitations
Currently, increasing the complexity of model inference to enhance system performance is a hot research topic Jaech et al. (2024); Guo et al. (2025). Our SpecFuse ensemble method not only significantly improves the response quality but also enhances the overall stability of the system by leveraging the strengths of multiple models during the inference stage. However, despite introducing a model exit mechanism to reduce the overall computational resource consumption during inference, model ensemble still requires more computational resources compared to using a single model. In particular, during the initial stage of inference, when all models participate, the demand for peak computational resources remains high. Looking ahead, we hope to combine SpecFuse with technologies such as model self-knowledge assessment, so that, before answering a question, the system can determine whether to include a model in the inference process based on the model’s confidence in its ability to answer the user’s question correctly. This approach would further optimize the efficiency of computational resource usage. We hope our work can inspire researchers and contribute to the ongoing progress in the field of model ensemble.
Appendix A The Motivation Of Model Exit Mechanism
The motivation for this approach stems from our observation that, when responding to a query, some models’ output segments rarely rank first. This indicates that these models are not well-suited for responding to the given query, making further computational investment in them inefficient. However, we cannot preemptively determine which models will underperform for a given query. Thus, we evaluate the contribution of each model to the query based on the cumulative quality scores of its previous candidate segments, as well as the frequency and position of its first-place rankings. We use the cumulative scores from previous rounds of each model as prior estimates of quality in subsequent rounds to determine whether a model should be exited. Since the number of rounds varies for different queries, a fixed threshold cannot be used for exit decisions. Therefore, we apply the softmax function to normalize the scores and set a temperature coefficient of $\sqrt{T}$ ($T$ being the current round). We choose $\sqrt{T}$ as the temperature coefficient because the number of output rounds rarely exceeds $100$. By using $\sqrt{T}$, we effectively limit the cumulative scores to under $10$, preventing extreme values from dominating. Additionally, we analyze the distribution of the best candidate segments. When these segments belong to only a few models, other models can be exited. By combining the temperature coefficient with the best segment distribution, the softmax scores more accurately reflect model performance, allowing underperforming models to exit promptly.
Appendix B Experiment Setups
B.1 Dataset Details
We evaluate all the models on six datasets that represent different core capabilities of LLMs, open-domain instruction-response (IR), commonsense, reasoning and instruction following.
@@ B.1 Dataset Details
  • Open-domain IR: We evaluate the model’s open-domain instruction-response capability using both English and Chinese human-computer interaction datasets. For the English dataset, we choose the Alpaca-gpt4 (Peng et al., 2023) and Dolly-15k (Conover et al., 2023) datasets for evaluation, both of which have inputs that consist of human instructions. We select these two datasets because their response sources differ: the Dolly-15k dataset features human-provided responses, while the Alpaca-GPT4 contains responses generated by the state-of-the-art GPT-4 (OpenAI et al., 2024) model, which provides neutral answers to each question and can refuse to answer inappropriate or harmful ones. Using both types of responses for scoring allows us to more thoroughly compare the advantages of our ensemble system. Additionally, due to the large size of these datasets, we randomly sample portions from each to create a new test set. From the Dolly-15k dataset, we randomly select 1,500 open-QA samples for testing, with 500 reserved for the development set. In the Alpaca-GPT4, after shuffling the data, we manually verify the correctness of GPT-4’s responses and select 2,000 validated samples, with 1,500 used for testing and 500 for validation. For the Chinese dataset, we utilize the Human-Value and Ruozb datasets from the COIG-CQIA (Bai et al., 2024) benchmark for testing. The instructions in these two datasets consist of human-posed questions, with answers provided either by humans or generated by GPT-4. The COIA authors manually review and filter the responses, retaining only the correct answers generated by GPT-4.
B.2 Evaluation Methods
We use a variety of metrics for different tasks, following the test scripts from the Openllm leaderboard. To assess the quality of human question-answering, we apply BARTScore (Bart-S) (Yuan et al., 2021), BERTScore (Bert-S) (Zhang et al., 2019), GPT4-Rank (GPT4-R) (OpenAI et al., 2024), BLEU (Papineni et al., 2002), and ROUGE (R-n) (Lin, 2004).
For multiple-choice tasks such as MMLU and ARC-C, we select the option with the highest likelihood to calculate accuracy (Acc). For the reasoning dataset GSM8K, we evaluate exact match (EM) accuracy. For IFEVAL, we rely on the evaluation files provided by the dataset creators (Zhou et al., 2023), testing under prompt-strict, instruction-strict, prompt-loose, and instruction-loose conditions.
A detailed description of the evaluation metrics for human question-answering is as follows:
@@ B.2 Evaluation Methods
  • The GPT4-Rank evaluation utilizes the GPT-4555The version we use is GPT-4-turbo, and the link is https://openai.com/index/gpt-4/ model to compare two different responses against a ground-truth response. The model will select the better of the two responses. For each test sample, we pair the responses generated by different models and have GPT-4 determine which one is superior. Since the MBR and PairRank methods do not generate new responses, we do not re-rank the responses they select from the base LLMs. Instead, we use the average rankings of the responses they select from the base LLMs to represent their GPT4-Rank. Once all comparisons are complete, we count the number of wins for each model. Based on these win counts, we rank the responses from the different models. The average ranking of each model across all data in the dataset is the value reported in our table. The evaluation instructions for GPT-4 are shown in Table 7.
B.3 Baselines
Since our approach has not undergone any additional training, we selecte several types of untrained baseline models for comparison with our method:
  • Larger LLMs: including Mixtral-8x7B-v0.1 (Jiang et al., 2023a), Qwen2-72B (Yang et al., 2024), and Llama-3-70B (AI@Meta, 2024).
  • PairRank: an English reward model introduced in the LLM-Blender (Jiang et al., 2023b), which compares candidate results generated by different LLMs and selects the best candidate as the ensemble output.
  • Minimum Bayes Risk (MBR) (Freitag et al., 2023): selects the answer with the highest lexical similarity to other candidate answers. In this paper, we use the SimCSE (Gao et al., 2022) model to calculate the similarity between candidate responses.
  • Generation Fusion (GF) (Jiang et al., 2023b): uses the outputs of other models as context, passing them to a new model, which generates a response based on this context.
  • Mixture-of-agents (MOA) Wang et al. (2024b): Multi-layer fusion is applied, where the outputs of all base models are concatenated and fed back into the models, with an Aggregator outputting the final result. In this paper, Mistral-7B-v0.3 is chosen as the aggressor, and the fusion process iterates three times, as in the original MOA paper.
  • Majority Voting Davani et al. (2022): each model provides a choice, and the final result is determined by the option with the most votes.
B.4 Implement details
As the methods in this paper are not trained, we only provide the parameter settings for inference. All models in this study are loaded with bfloat16 precision for inference and use the following generation parameters: $do\_sample=True$, $temperature=0.6$, and $top\_p=0.9$. All of our experiments are conducted on A100 GPUs, and set the maximum length of candidate segments to $10$. All the experimental results in this paper are obtained by testing three times and averaging the values.
Model Bert-S RougeL Avg. Models/Reply
SpecFuse 72.8354 28.3507 2.4168
[…]
     w/o Q Score (n=5) 71.7152 25.1712 1.9366
     w/o Q Score (n=10) 72.1122 27.3213 3.2562
     w/o Q Score (n=15) 72.8402 28.4322 4.3280
SpecFuse($\tau=1$) 71.7362 25.1196 2.0268
SpecFuse ($\tau=\sqrt{T}$) 72.8105 28.2746 4.1996
Table 6: Ablation study of the Model Exit mechanism. “w/o Q Score (n=$5$)” means that the cumulative quality score is not applied, and models that fail to generate the best segment within the first $5$ segments (n=$5$) are excluded. “w/o entropy” indicates that entropy is not used to dynamically adjust the softmax temperature coefficient. SpecFuse($\tau=1$) indicates that the Model Exit Mechanism is used but without dynamic temperature scaling, with the temperature fixed at $1$, and ($\tau=\sqrt{T}$) indicates the temperature is fixed at $\sqrt{T}$, where $T$ refers to the current iteration of SpecFuse. “Avg. Models/Reply” refers to the average number of models invoked by our framework to respond each question.
@@ 2. B is better
Appendix C Experimental Analysis
C.1 Analysis of Model Exit Mechanism
To more intuitively observe whether the Model Exit Mechanism removes unimportant models, we designed a comparison of the proportion of best segment outputs from each model, both with and without the mechanism applied. As shown in Figure 5, the proportion of best candidate generations for each LLM changes only slightly with or without Model Exit Mechanism, indicating that the mechanism accurately eliminates LLMs with low selection probability.
C.2 Ablation study of the Model Exit mechanism
From the Table 6, it can be observed that methods not using the cumulative quality score significantly reduce the number of model calls when n=5. However, performance also drops substantially because the best candidate does not fully reflect the quality of a model’s output. Some models may not be selected in the initial rounds but generate candidates with scores only slightly lower than the best. These models are likely to produce the best candidate in later rounds. As n increases, performance metrics improve noticeably, but this comes at the cost of a significant rise in the number of model calls, leading to increased computational resource requirements. Therefore, choosing an appropriate n is highly challenging. Moreover, the table also shows that without using entropy, the reduction in computational resources is not significant. This is because, when the cumulative scores are not substantially different, the normalized scores show minimal distinction, resulting in models with little contribution not being exited in a timely manner.
We also conduct an ablation experiment with a fixed temperature coefficient, as shown in Table 5, where SpecFuse ($\tau=1$) results in the fewest model invocations but suffers from significant performance loss. This occurs because, as iterations increase, the cumulative model scores grow larger, causing the softmax function to produce a sharper distribution. Consequently, models with slightly lower scores are prematurely eliminated from the response, which negatively impacts the overall performance of the ensemble. SpecFuse ($\tau=\sqrt{T}$) sets the temperature coefficient to $\sqrt{T}$, which makes the softmax function overly smooth. As a result, it takes many iterations to accumulate substantial score differences between models before the lower-scoring models exit the current response, leading to delayed exits and an excessive number of model invocations overall. In contrast, the Model Exit mechanism, with dynamic temperature scaling, adjusts the temperature coefficient based on the distribution of the best candidate from previous rounds, ensuring timely model exits.
C.3 Time Consumption Experiment Setup
The base models in the ensemble are Qwen2-7B, GLM-4-9B, Llama-3-8B, and Mistral-7B-v0.3. GenFuse uses GLM-4 as the fusion model, while MOA also employs GLM-4 as the aggregator. The experiments are conducted on a setup with 4*A100-80G GPUs, , and all methods are tested under the same conditions without the use of inference acceleration tools. To test, we use the prompt: "Write a 10,000-word essay on the history of the Roman Empire" to generate long-form text. We set the maximum lengths to $100$, $500$, $1000$, $1500$, and $2000$ to measure the time required by each method to generate outputs of these lengths. Each result represents the average value calculated from $20$ tests.
Appendix D Case Study